If x > y > 0, x2 + y2= 13 and x y = 6. Then y = ?
 

Grade:10

1 Answers

Sunil Raikwar
9 years ago
Hello student, please check the solution of your question given below

x^2+y^2+2xy=13+2*6
(x+y)^2=25
x+y=5.......................1

x^2+y^2-2xy=13-2*6
(x-y)^2=1
x-y=1..............................2
solving equation 1 & 2
we get x=3 & y=2